Shift report 20010907 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Fri Sep 07 2001 - 00:02:40 EDT

  • Next message: Apache: "Shift report 20010907 00:00-08:00"

    ------------------------------------------------------------
    Date: Fri Sep  7 00:02:40 2001, Shift: 16:00-24:00
    Supervisor: A. Jippa, with: K. Hagel
    Log book#8 pages 88 - 89
    
    Shift Summary:
    ------------------------------------------------------------
    No beam during entire shift.
    Machine problems with Quench Link Interlock again.
    Turned on Magnets again:
    D1 676 Pol A
    D2 552 Pol A
    D3 585 Pol A
    D4 538 Pol A
    D5 500 Pol A
    
    FV added a readout of 3 words to trigger word in order to
    test vertex trigger.
    KH added extra 3 words to BrEventHeader, tested, submitted
    to CVS and updated operator account.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Fri Sep 07 2001 - 00:03:08 EDT